# Generalizable Domain Adaptation for Sim-and-Real Policy Co-Training

### v2

@@ Abstract
Behavior cloning has shown promise for robot manipulation, but real-world demonstrations are costly to acquire at scale. While simulated data offers a scalable alternative, particularly with advances in automated demonstration generation, transferring policies to the real world is hampered by various simulation and real domain gaps. In this work, we propose a unified sim-and-real co-training framework for learning generalizable manipulation policies that primarily leverages simulation and only requires a few real-world demonstrations. Central to our approach is learning a domain-invariant, task-relevant feature space. Our key insight is that aligning the joint distributions of observations and their corresponding actions across domains provides a richer signal than aligning observations (marginals) alone. We achieve this by embedding an Optimal Transport (OT)-inspired loss within the co-training framework, and extend this to an Unbalanced OT framework to handle the imbalance between abundant simulation data and limited real-world examples. We validate our method on challenging manipulation tasks, showing it can leverage abundant simulation data to achieve up to a 30% improvement in the real-world success rate and even generalize to scenarios seen only in simulation.
@@ 1 Introduction
Behavior cloning pomerleau1988alvinn is a promising approach for acquiring robot manipulation skills directly in the real world, due to its simplicity and effectiveness in mimicking expert demonstrations robomimic2021 ; florence2022implicit . However, achieving robust and generalizable performance requires collecting large-scale datasets khazatsky2024droid ; open_x_embodiment_rt_x_2023 across diverse environments, object configurations, and tasks. This data collection process is labor-intensive, time-consuming, and costly, posing significant challenges to scalability in real-world applications.
Recently, with rapid advancements in physics simulators Genesis ; Xiang_2020_SAPIEN , procedural scene generation raistrick2024infinigen ; deitke2022️ , and motion synthesis techniques mandlekar2023mimicgen ; cheng2023nod , there has been growing interest in leveraging simulation as an alternative source of training data. These simulation-based approaches enable scalable and controllable data generation, allowing for diverse and abundant supervision at a fraction of the real-world cost. However, transferring policies trained in simulation to the physical world remains a non-trivial challenge due to sim-to-real gap—the discrepancies between the simulated and real-world environments that a policy encounters during execution. These differences can manifest in various forms, such as variations in visual appearance, sensor noise, and action dynamics andrychowicz2020learning ; tobin2017domain . In particular, learning visuomotor control policies that remain robust under changing perceptual conditions during real-world deployment continues to be an open area of research.
@@
Common strategies to bridge this domain gap include domain randomization andrychowicz2020learning ; tobin2017domain and data augmentation hansen2020self ; yarats2021mastering , though these often require careful tuning. Domain adaptation (DA) techniques aim to explicitly align distributions, either at pixel bousmalis2017unsupervised ; james2019sim or feature levels tzeng2014deep ; long2015learning ; zhao2019learning . However, many feature-level methods align only marginal observation distributions (e.g., MMD tzeng2014deep ; long2015learning ), which can be insufficient for fine-grained manipulation alignment as it may not preserve action-relevant relationships across domains. More recently, sim-and-real co-training—simply training a single policy on mixed data from both domains wei2025empirical ; maddukuri2025sim —has shown surprising effectiveness. We argue that while beneficial for data diversity, such co-training approaches typically lack explicit constraints for feature space alignment across domains, potentially hindering optimal transfer and generalization because they don’t enforce a consistent mapping of task-relevant structures.
We present a unified sim-and-real co-training framework that explicitly learns a shared latent space where observations from simulation and the real world are aligned and preserve action-relevant information. Our key insight is that aligning the joint distributions of observations and their corresponding actions or task-relevant states across domains provides a direct signal for learning transferable features. Concretely, we leverage Optimal Transport (OT) courty2016optimal as an alignment objective to learn representations where the geometric relationships crucial for action prediction are consistent, irrespective of whether the input comes from simulation or the real world. Further more, to robustly handle the data imbalance in co-training with abundant simulation data and limited real-world data, we further extend to an Unbalanced OT (UOT) formulation fatras2021unbalanced ; chizat2018scaling and develop a temporally-aware sampling strategy to improve domain alignment learning in a mini-batch OT setting.
@@ 2 Related Work
Behavior Cloning for Robot Manipulation. Behavior cloning (BC) trains policies to map observations to actions by imitating expert demonstrations robomimic2021 ; zhao2023learning ; chi2023diffusion ; Ze2024DP3 , offering an effective path to human-like manipulation skills. Generalization heavily depends on dataset diversity. While some efforts focus on large-scale real-world data collection open_x_embodiment_rt_x_2023 ; khazatsky2024droid or more efficient collection techniques wu2024gello ; chi2024universal ; kareer2024egomimic , this remains costly and time-consuming. Physical simulators provide a low-cost alternative, with automatic motion synthesis leveraging privileged information to generate large-scale simulated demonstrations mandlekar2023mimicgen ; jiang2024dexmimicgen ; cheng2024nodtamp . Our work combines abundant simulated data with few real-world demonstrations to train robust BC policies.
@@
Sim-to-real Transfer and Co-training. Policies trained solely in simulation often underperform in the real world due to the sim-to-real gap—discrepancies in visual appearance and dynamics. For quasi-static manipulation, the visual domain gap is typically the primary bottleneck. Domain randomization exposes policies to varied simulated visual conditions to build robustness to real-world variability tobin2017domain ; james2017transferring ; yuan2024learning . However, its success depends on how well randomized parameters cover true real-world distributions, often requiring manual tuning. Domain adaptation (DA) explicitly aligns source (simulation) and target (real) domains farahani2021brief . Pixel-level DA uses image translation to make simulated images resemble real ones bousmalis2017unsupervised ; james2019sim ; ho2021retinagan . Feature-level DA, which is often more scalable for end-to-end learning, focuses on learning domain-invariant representations long2015learning ; zhao2019learning ; tzeng2014deep ; raychaudhuri2021cross ; kim2020domain . Sim-and-real cotraining, where a policy is jointly trained on mixed data wei2025empirical ; maddukuri2025sim , offers a simple and effective alternative. While cotraining enhances generalization through data diversity, it typically lacks explicit constraints to align learned feature spaces across domains. We build upon cotraining by incorporating feature-level domain adaptation via Optimal Transport to promote latent space alignment, thereby improving real-world policy performance.
@@
Optimal Transport for Domain Adaptation. Optimal Transport (OT) offers a principled framework for aligning distributions, widely adopted for domain adaptation courty2016optimal ; courty2014domain ; perrot2016mapping ; redko2017theoretical ; courty2017joint ; nguyen2024dude . Traditional OT methods compute a transport plan between source and target samples, then train a new model on the transported source. Most relevant to us is DeepJDOT damodaran2018deepjdot , which builds on JDOT courty2017joint to align joint distributions of features and (pseudo) labels for unsupervised domain adaptation, where target labels are unavailable. Our work builds on these principles for sim-and-real co-training imitation. Unlike unsupervised DA, we leverage available action or state labels from limited real-world demonstrations as “soft” supervision to guide a more task-relevant alignment of joint observation-label distributions across domains. To robustly handle the inherent data imbalance between abundant simulation and scarce real data, we incorporate an Unbalanced OT (UOT) loss fatras2021unbalanced into our co-training framework and develop a temporally-aware sampling strategy to improve mini-batch UOT training.
@@ 3.1 Problem Setting: Sim-and-Real Policy Co-Training
We address the challenge of learning robust real-world robotic manipulation policies $\pi$. Our approach minimizes the need for extensive real-world data collection by primarily leveraging abundant simulation data alongside a small set of real-world demonstrations. This is framed as a sim-and-real co-training problem maddukuri2025sim ; matas2018sim , where a single policy is trained on data from both domains. Specifically, we consider a source domain (simulation, denoted srcsrc) and a target domain (real-world, denoted tgttgt). We model the domains as Partially Observable Markov Decision Processes (POMDPs) that share an underlying, generally unobserved, state space $\mathcal{S}$ and an action space $\mathcal{A}$. The policy receives observations comprising high-dimensional visual input $o\in\mathcal{O}$ (e.g., RGB images, 3D point clouds) generated by the emission function $E:\mathcal{S}\mapsto\mathcal{O}$, together with low-dimensional proprioceptive information $x\in\mathcal{X}$ (e.g., robot joint angles, end-effector pose).
@@
Data Sources and Data Imbalance. In the source (simulation) domain, we leverage the ability to automatically generate a large dataset of $N_{src}$ trajectories, $D_{src}=\{(o^{i}_{src},x^{i}_{src},a^{i}_{src})\}_{i=1}^{N_{src}}$. Specifically, we leverage an automated demonstration generation tool MimicGen mandlekar2023mimicgen , which utilizes privileged information available in simulation to create a diverse set of experiences covering a broad underlying state space, $S_{src}$. MimicGen populates synthetic demonstrations based on a handful of human demonstrations, which ensures that the generated data is behaviorally consistent with human demonstrations. We collect a limited number of $N_{tgt}$ demonstrations, $D_{tgt}=\{(o^{j}_{tgt},x^{j}_{tgt},a^{j}_{tgt})\}_{j=1}^{N_{tgt}}$ (where $N_{src}\gg N_{tgt}$), typically through human teleoperation. These real-world demonstrations will naturally cover a much smaller and potentially distinct subset of states, $S_{tgt}$. This difference in data coverage leads to the challenge of partial data overlap. While we assume there is a region of states common to both $S_{src}$ and $S_{tgt}$ where direct alignment is possible, a significant portion of $S_{src}$ (our rich simulated data) will not have corresponding real-world demonstrations in $S_{tgt}$. Conversely, $S_{tgt}$ might contain details specific to real world (e.g., demonstration behaviors). Effectively leveraging the entirety of $D_{src}$ for real-world performance, especially for states outside the direct sim-real overlap in demonstrations, is the problem we address.
@@
Objective: Generalizable Domain Adaptation. Our goal is to learn a single, generalizable policy $\pi_{\theta}(a|z,x)$ and an observation encoder $f_{\phi}:\mathcal{O}_{src}\cup\mathcal{O}_{tgt}\rightarrow\mathcal{Z}$. This encoder maps high-dimensional visual observations $o$ from both source and target domains to a shared latent space $\mathcal{Z}$. The primary objective is to achieve high policy performance in the target (real-world) domain, especially in scenarios not explicitly covered by the limited real-world demonstrations $D_{tgt}$. This entails two objectives. First, for states within the overlapping regions of $S_{src}$ and $S_{tgt}$, we aim to learn high-quality embeddings $z=f_{\phi}(o)$ that are well-aligned across domains, such that $f_{\phi}(o_{src})\approx f_{\phi}(o_{tgt})$ in corresponding states, facilitating effective policy learning. This is also the assumption of most co-training methods wei2025empirical ; maddukuri2025sim . Second, for states covered in $S_{src}$ but not in $S_{tgt}$, the encoder $f_{\phi}$ must produce embeddings $f_{\phi}(o_{tgt})$ for novel target observations that are consistent with the embeddings of their simulated counterparts $f_{\phi}(o_{src})$. This requires the learned representations to capture domain-invariant, task-relevant features, enabling policy to generalize to target (real-world) scenarios for which training data is only present in the source (simulated) domain.
@@ 4.1 Optimal Transport for Action-Aware Feature Alignment
While standard co-training methods wei2025empirical ; maddukuri2025sim offer implicit feature alignment, and marginal distribution matching (e.g., Maximum Mean Discrepancy tzeng2014deep ) can overlook fine-grained correspondences, we seek a more structured approach. To learn domain-invariant visual features $z=f_{\phi}(o)$ that are also predictive of actions $a\in\mathcal{A}$, we propose to align the joint distributions $P_{src}(f_{\phi}(o_{src}),a_{src})$ and $P_{tgt}(f_{\phi}(o_{tgt}),a_{tgt})$ using Optimal Transport. This encourages the encoder $f_{\phi}$ to learn representations where the geometric relationships between (visual feature, action) pairs are preserved across domains. By minimizing an OT-based loss, the encoder $f_{\phi}$ is trained to shape the embedding space $\mathcal{Z}$ such that structures relevant to action prediction are consistent between simulation and the real world.
@@
Minimizing this objective using an iterative algorithm like Sinkhorn cuturi2013sinkhorn creates a bi-level optimization. In the inner loop, for a fixed $f_{\phi}$, an approximately optimal transport plan $\Pi$ is computed. In the outer loop, $f_{\phi}$ is updated to reduce the cost incurred by this plan. This process effectively trains the encoder $f_{\phi}$ to produce embeddings $z$ that make the source and target joint distributions $(z,a)$ less costly to align. A key advantage of OT is its ability to preserve geometric structures; by guiding alignment with action similarity (via $d_{\mathcal{A}}$), we shape the embedding function $f_{\phi}$ to cluster visual observations that lead to similar actions, irrespective of their domain of origin.

### v3

@@ Abstract
Behavior cloning has shown promise for robot manipulation, but real-world demonstrations are costly to acquire at scale. While simulated data offers a scalable alternative, particularly with advances in automated demonstration generation, transferring policies to the real world is hampered by various simulation and real domain gaps. In this work, we propose a unified sim-and-real co-training framework for learning generalizable manipulation policies that primarily leverages simulation and only requires a few real-world demonstrations. Central to our approach is learning a domain-invariant, task-relevant feature space. Our key insight is that aligning the joint distributions of observations and their corresponding actions across domains provides a richer signal than aligning observations (marginals) alone. We achieve this by embedding an Optimal Transport (OT)-inspired loss within the co-training framework, and extend this to an Unbalanced OT framework to handle the imbalance between abundant simulation data and limited real-world examples. We validate our method on challenging manipulation tasks, showing it can leverage abundant simulation data to achieve up to a 30% improvement in the real-world success rate and even generalize to scenarios seen only in simulation. Project webpage: https://ot-sim2real.github.io/.
@@ 1 Introduction
Behavior cloning Pomerleau (1988) is a promising approach for acquiring robot manipulation skills directly in the real world, due to its simplicity and effectiveness in mimicking expert demonstrations Mandlekar et al. (2021); Florence et al. (2022). However, achieving robust and generalizable performance requires collecting large-scale datasets Khazatsky et al. (2024); Collaboration et al. (2023) across diverse environments, object configurations, and tasks. This data collection process is labor-intensive, time-consuming, and costly, posing significant challenges to scalability in real-world applications.
Recently, with rapid advancements in physics simulators Authors (2024); Xiang et al. (2020), procedural scene generation Raistrick et al. (2024); deitke2022️, and motion synthesis techniques Mandlekar et al. (2023); Cheng et al. (2023), there has been growing interest in leveraging simulation as an alternative source of training data. These simulation-based approaches enable scalable and controllable data generation, allowing for diverse and abundant supervision at a fraction of the real-world cost. However, transferring policies trained in simulation to the physical world remains a non-trivial challenge due to sim-to-real gap—the discrepancies between the simulated and real-world environments that a policy encounters during execution. These differences can manifest in various forms, such as variations in visual appearance, sensor noise, and action dynamics Andrychowicz et al. (2020); Tobin et al. (2017). In particular, learning visuomotor control policies that remain robust under changing perceptual conditions during real-world deployment continues to be an open area of research.
@@
Common strategies to bridge this domain gap include domain randomization Andrychowicz et al. (2020); Tobin et al. (2017) and data augmentation Hansen et al. (2020); Yarats et al. (2021), though these often require careful tuning. Domain adaptation (DA) techniques aim to explicitly align distributions, either at pixel Bousmalis et al. (2017); James et al. (2019) or feature levels Tzeng et al. (2014); Long et al. (2015); Zhao et al. (2019). However, many feature-level methods align only marginal observation distributions (e.g., MMD Tzeng et al. (2014); Long et al. (2015)), which can be insufficient for fine-grained manipulation alignment as it may not preserve action-relevant relationships across domains. More recently, sim-and-real co-training—simply training a single policy on mixed data from both domains Wei et al. (2025); Maddukuri et al. (2025)—has shown surprising effectiveness. We argue that while beneficial for data diversity, such co-training approaches typically lack explicit constraints for feature space alignment across domains, potentially hindering optimal transfer and generalization because they don’t enforce a consistent mapping of task-relevant structures.
We present a unified sim-and-real co-training framework that explicitly learns a shared latent space where observations from simulation and the real world are aligned and preserve action-relevant information. Our key insight is that aligning the joint distributions of observations and their corresponding actions or task-relevant states across domains provides a direct signal for learning transferable features. Concretely, we leverage Optimal Transport (OT) Courty et al. (2016) as an alignment objective to learn representations where the geometric relationships crucial for action prediction are consistent, irrespective of whether the input comes from simulation or the real world. Further more, to robustly handle the data imbalance in co-training with abundant simulation data and limited real-world data, we further extend to an Unbalanced OT (UOT) formulation Fatras et al. (2021); Chizat et al. (2018) and develop a temporally-aware sampling strategy to improve domain alignment learning in a mini-batch OT setting.
@@ 2 Related Work
Behavior Cloning for Robot Manipulation. Behavior cloning (BC) trains policies to map observations to actions by imitating expert demonstrations Mandlekar et al. (2021); Zhao et al. (2023); Chi et al. (2023b); Ze et al. (2024), offering an effective path to human-like manipulation skills. Generalization heavily depends on dataset diversity. While some efforts focus on large-scale real-world data collection Collaboration et al. (2023); Khazatsky et al. (2024) or more efficient collection techniques Wu et al. (2024); Chi et al. (2024); Kareer et al. (2024), this remains costly and time-consuming. Physical simulators provide a low-cost alternative, with automatic motion synthesis leveraging privileged information to generate large-scale simulated demonstrations Mandlekar et al. (2023); Jiang et al. (2024); Cheng et al. (2024). Our work combines abundant simulated data with few real-world demonstrations to train robust BC policies.
@@
Sim-to-real Transfer and Co-training. Policies trained solely in simulation often underperform in the real world due to the sim-to-real gap—discrepancies in visual appearance and dynamics. For quasi-static manipulation, the visual domain gap is typically the primary bottleneck. Domain randomization exposes policies to varied simulated visual conditions to build robustness to real-world variability Tobin et al. (2017); James et al. (2017); Yuan et al. (2024). However, its success depends on how well randomized parameters cover true real-world distributions, often requiring manual tuning. Domain adaptation (DA) explicitly aligns source (simulation) and target (real) domains Farahani et al. (2021); Dan et al. (2025). Pixel-level DA uses image translation to make simulated images resemble real ones Bousmalis et al. (2017); James et al. (2019); Ho et al. (2021). Feature-level DA, which is often more scalable for end-to-end learning, focuses on learning domain-invariant representations Long et al. (2015); Zhao et al. (2019); Tzeng et al. (2014); Raychaudhuri et al. (2021); Kim et al. (2020). Sim-and-real co-training, where a policy is jointly trained on mixed data Wei et al. (2025); Maddukuri et al. (2025), offers a simple and effective alternative. While co-training enhances generalization through data diversity, it typically lacks explicit constraints to align learned feature spaces across domains. We build upon co-training by incorporating feature-level domain adaptation via Optimal Transport to promote latent space alignment, thereby improving real-world policy performance.
@@
Optimal Transport for Domain Adaptation. Optimal Transport (OT) offers a principled framework for aligning distributions, widely adopted for domain adaptation Courty et al. (2016, 2014); Perrot et al. (2016); Redko et al. (2017); Courty et al. (2017); Nguyen et al. (2024); Kedia et al. (2025). Traditional OT methods compute a transport plan between source and target samples, then train a new model on the transported source. Most relevant to us is DeepJDOT Damodaran et al. (2018), which builds on JDOT Courty et al. (2017) to align joint distributions of features and (pseudo) labels for unsupervised domain adaptation, where target labels are unavailable. Our work builds on these principles for sim-and-real co-training imitation. Unlike unsupervised DA, we leverage available action or state labels from limited real-world demonstrations as “soft” supervision to guide a more task-relevant alignment of joint observation-label distributions across domains. To robustly handle the inherent data imbalance between abundant simulation and scarce real data, we incorporate an Unbalanced OT (UOT) loss Fatras et al. (2021) into our co-training framework and develop a temporally-aware sampling strategy to improve mini-batch UOT training.
@@ 3.1 Problem Setting: Sim-and-Real Policy Co-Training
We address the challenge of learning robust real-world robotic manipulation policies $\pi$. Our approach minimizes the need for extensive real-world data collection by primarily leveraging abundant simulation data alongside a small set of real-world demonstrations. This is framed as a sim-and-real co-training problem Maddukuri et al. (2025); Matas et al. (2018), where a single policy is trained on data from both domains. Specifically, we consider a source domain (simulation, denoted srcsrc) and a target domain (real-world, denoted tgttgt). We model the domains as Partially Observable Markov Decision Processes (POMDPs) that share an underlying, generally unobserved, state space $\mathcal{S}$ and an action space $\mathcal{A}$. The policy receives observations comprising high-dimensional visual input $o\in\mathcal{O}$ (e.g., RGB images, 3D point clouds) generated by the emission function $E:\mathcal{S}\mapsto\mathcal{O}$, together with low-dimensional proprioceptive information $x\in\mathcal{X}$ (e.g., robot joint angles, end-effector pose).
@@
Data Sources and Data Imbalance. In the source (simulation) domain, we leverage the ability to automatically generate a large dataset of $N_{src}$ trajectories, $D_{src}=\{(o^{i}_{src},x^{i}_{src},a^{i}_{src})\}_{i=1}^{N_{src}}$. Specifically, we leverage an automated demonstration generation tool MimicGen Mandlekar et al. (2023), which utilizes privileged information available in simulation to create a diverse set of experiences covering a broad underlying state space, $S_{src}$. MimicGen populates synthetic demonstrations based on a handful of human demonstrations, which ensures that the generated data is behaviorally consistent with human demonstrations. We collect a limited number of $N_{tgt}$ demonstrations, $D_{tgt}=\{(o^{j}_{tgt},x^{j}_{tgt},a^{j}_{tgt})\}_{j=1}^{N_{tgt}}$ (where $N_{src}\gg N_{tgt}$), typically through human teleoperation. These real-world demonstrations will naturally cover a much smaller and potentially distinct subset of states, $S_{tgt}$. This difference in data coverage leads to the challenge of partial data overlap. While we assume there is a region of states common to both $S_{src}$ and $S_{tgt}$ where direct alignment is possible, a significant portion of $S_{src}$ (our rich simulated data) will not have corresponding real-world demonstrations in $S_{tgt}$. Conversely, $S_{tgt}$ might contain details specific to real world (e.g., demonstration behaviors). Effectively leveraging the entirety of $D_{src}$ for real-world performance, especially for states outside the direct sim-real overlap in demonstrations, is the problem we address.
@@
Objective: Generalizable Domain Adaptation. Our goal is to learn a single, generalizable policy $\pi_{\theta}(a|z,x)$ and an observation encoder $f_{\phi}:\mathcal{O}_{src}\cup\mathcal{O}_{tgt}\rightarrow\mathcal{Z}$. This encoder maps high-dimensional visual observations $o$ from both source and target domains to a shared latent space $\mathcal{Z}$. The primary objective is to achieve high policy performance in the target (real-world) domain, especially in scenarios not explicitly covered by the limited real-world demonstrations $D_{tgt}$. This entails two objectives. First, for states within the overlapping regions of $S_{src}$ and $S_{tgt}$, we aim to learn high-quality embeddings $z=f_{\phi}(o)$ that are well-aligned across domains, such that $f_{\phi}(o_{src})\approx f_{\phi}(o_{tgt})$ in corresponding states, facilitating effective policy learning. This is also the assumption of most co-training methods Wei et al. (2025); Maddukuri et al. (2025). Second, for states covered in $S_{src}$ but not in $S_{tgt}$, the encoder $f_{\phi}$ must produce embeddings $f_{\phi}(o_{tgt})$ for novel target observations that are consistent with the embeddings of their simulated counterparts $f_{\phi}(o_{src})$. This requires the learned representations to capture domain-invariant, task-relevant features, enabling policy to generalize to target (real-world) scenarios for which training data is only present in the source (simulated) domain.
@@ 4.1 Optimal Transport for Action-Aware Feature Alignment
While standard co-training methods Wei et al. (2025); Maddukuri et al. (2025) offer implicit feature alignment, and marginal distribution matching (e.g., Maximum Mean Discrepancy Tzeng et al. (2014)) can overlook fine-grained correspondences, we seek a more structured approach. To learn domain-invariant visual features $z=f_{\phi}(o)$ that are also predictive of actions $a\in\mathcal{A}$, we propose to align the joint distributions $P_{src}(f_{\phi}(o_{src}),a_{src})$ and $P_{tgt}(f_{\phi}(o_{tgt}),a_{tgt})$ using Optimal Transport. This encourages the encoder $f_{\phi}$ to learn representations where the geometric relationships between (visual feature, action) pairs are preserved across domains. By minimizing an OT-based loss, the encoder $f_{\phi}$ is trained to shape the embedding space $\mathcal{Z}$ such that structures relevant to action prediction are consistent between simulation and the real world.
@@
Minimizing this objective using an iterative algorithm like Sinkhorn Cuturi (2013) creates a bi-level optimization. In the inner loop, for a fixed $f_{\phi}$, an approximately optimal transport plan $\Pi$ is computed. In the outer loop, $f_{\phi}$ is updated to reduce the cost incurred by this plan. This process effectively trains the encoder $f_{\phi}$ to produce embeddings $z$ that make the source and target joint distributions $(z,a)$ less costly to align. A key advantage of OT is its ability to preserve geometric structures; by guiding alignment with action similarity (via $d_{\mathcal{A}}$), we shape the embedding function $f_{\phi}$ to cluster visual observations that lead to similar actions, irrespective of their domain of origin.